OpenStudy (anonymous):

Determine the sum or difference in clock 12 arithmetic. (6 + 1) + 12 A. 19 B. 8 C. 7 D. 3

OpenStudy (andriod09):

clock 12 arithmatic?

OpenStudy (anonymous):

??

OpenStudy (kymber):

C, 7! :]

OpenStudy (anonymous):

thanks

OpenStudy (andriod09):

6+1=7 7+12= 15 15-12=3

OpenStudy (kymber):

6 + 1 is 7. 12 hours later than 7 is 7. I must be doing this wrong, how could it be 3?

OpenStudy (andriod09):

7+12=15 15-12=3

jimthompson5910 (jim_thompson5910):

(6+1) + 12 7 + 12 19 .... oops, made a mistake here.... 7 So (6+1) + 12 = 7 (mod 12)

OpenStudy (kymber):